OpenStudy (anonymous):

Earl is graduating from college in one year, but he will need a loan in the amount of $3,650 for his last two semesters. He may either receive an unsubsidized Stafford Loan with an interest rate of 6.8%, compounded monthly, or his parents may get a PLUS Loan with an interest rate of 7.8%, compounded monthly. The Stafford Loan has a grace period of six months from the time of graduation. Which loan will have a higher balance and by how much at the time of repayment?

OpenStudy (anonymous):

a) The Standford Loan has a higher balance by $95.69 b) The PLUS Loan has a higher balance by $60.69 c) The PLUS Load has a higher balance by $39.02 d) The Standford Loan has a higher balance by $246.11

jimthompson5910 (jim_thompson5910):

Did you get anywhere with this one? Did you get started at all?

OpenStudy (anonymous):

No. I'm just really lost.

jimthompson5910 (jim_thompson5910):

Hint: Use the formula A = P(1 + r/n)^(nt) and keep the grace period of 6 months in mind (ie you'll have to add 6 months to the time t for the stafford loan)

OpenStudy (anonymous):

Thank you. I feel silly because you posted the same formula on my other question, and I could've just used it for this one. But thanks

jimthompson5910 (jim_thompson5910):

Let's compute the balance for the stafford loan A = P(1 + r/n)^(nt) A = 3650(1 + 0.068/12)^(12*18/12) A = ??? Note: use t = 18/12 because t is in years and instead of using t = 12/12 or t = 1 year, you have to add on the 6 month grace period where interest is still accrued.

jimthompson5910 (jim_thompson5910):

So what do you get

OpenStudy (anonymous):

3818.79

jimthompson5910 (jim_thompson5910):

I'm not getting that

jimthompson5910 (jim_thompson5910):

I'm getting 4040.79

OpenStudy (anonymous):

Hmm...maybe I typed it in incorrectly. I'll try again

OpenStudy (anonymous):

ok that's what I got. I put in 12/18 instead of 18/12

jimthompson5910 (jim_thompson5910):

ah i gotcha

jimthompson5910 (jim_thompson5910):

now compute the balance for the PLUS loan

jimthompson5910 (jim_thompson5910):

A = P(1 + r/n)^(nt) A = 3650(1 + 0.078/12)^(12*12/12) now t = 12/12 or just t = 1 A = ???

OpenStudy (anonymous):

I got 3945.10

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

now subtract the two

OpenStudy (anonymous):

95.69

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

thanks, again! :)

jimthompson5910 (jim_thompson5910):

you're welcome